OpenStudy (anonymous):

Choose the procedure you would use to solve the equation 3x - 3 = 35 for x. Add three to both sides, and then divide both sides by three. Add three to both sides, and then multiply both sides by three. Divide both sides by three, and then subtract three from both sides. Multiply both sides by three, and then add three to both sides.

OpenStudy (anonymous):

Add three to both sides, and then divide both sides by three.

OpenStudy (anonymous):

just like the last one: 3x - 3 = 35 +3 +3 3x = 38 x = 38/3